# The 6th Affective Behavior Analysis in-the-wild (ABAW) Competition

### 29 Feb 2024 v1

@@ 1 Introduction
For this Challenge, an augmented version of the Aff-Wild2 [105, 44, 47, 37, 51, 40, 39, 54, 30, 31, 48, 59] is used. This corpora is audiovisual (A/V), in-the-wild and in total consists of 594 videos of around 3M frames of 584 subjects.
@@
Only uni-task solutions will be accepted for this Challenge. Teams are allowed to use any -publicly or not- available pre-trained model (as long as it has not been pre-trained on the utilized in this Challenge corpora, Aff-Wild2). The pre-trained model can be pre-trained on any task (e.g., VA estimation, Expression Recognition, AU detection, Face Recognition). However when the teams are refining the model and developing the methodology, they should not use any other annotations (VA or AUs): the methodology should be purely uni-task, using only the Expr annotations. This means that teams are allowed to use other databases’ Expr annotations, or generated/synthetic data (e.g. the data provided in the ECCV 2022 run of the ABAW Challenge [29]), or any affine transformations, or in general data augmentation techniques (e.g., [78, 60]) for increasing the size of the training dataset.
@@
Only uni-task solutions will be accepted for this Challenge. Teams are allowed to use any -publicly or not- available pre-trained model (as long as it has not been pre-trained on the utilized in this Challenge corpora, Aff-Wild2). The pre-trained model can be pre-trained on any task (e.g., VA estimation, Expression Classification, AU detection, Face Recognition). However when the teams are refining the model and developing the methodology, they should not use any other annotations (VA or Expr): the methodology should be purely uni-task, using only the AU annotations. This means that teams are allowed to use other databases’ AU annotations, or generated/synthetic data, or any affine transformations, or in general data augmentation techniques (e.g., [78, 60]) for increasing the size of the training dataset.
@@
The sixth ABAW Competition, which is part of the respective Workshop to be held in conjunction with the IEEE Computer Vision and Pattern Recognition Conference (CVPR) 2024 is a continuation of the successful series of ABAW Competitions held in conjunction with IEEE CVPR 2023, ECCV 2022, IEEE CVPR 2022, ICCV 2021, IEEE FG 2020 and IEEE CVPR 2017, with the participation of many teams coming from both academia and industry, from all across the world [113, 108, 97, 84, 91, 96, 101, 115, 58, 114, 113, 84, 100, 104, 17, 73, 71, 115, 85, 28, 10, 65, 112, 12, 6, 4, 11, 110, 90, 92, 106, 98, 25, 1, 75, 62, 14, 13, 102, 68, 15, 69, 76, 23, 19, 9, 80, 70, 107, 74, 82, 26, 79, 111, 103, 27, 77, 99, 21, 95, 20, 89, 94, 24, 8, 18, 86, 5, 16, 66, 93, 72, 83, 67, 109, 63, 22, 64, 56, 3, 57, 55, 2, 53, 49, 52, 61, 81, 42, 38, 43, 46, 35, 33, 34, 36, 87, 88, 45, 50, 41].
@@ 2.5 Emotional Mimicry Intensity Estimation Challenge
For this Challenge, the multimodal Hume-Vidmimic2 dataset is used which consists of more than 15,000 videos totaling over 25 hours. In this dataset, every participant was tasked with imitating a ’seed’ video that showcased an individual displaying a particular emotion. Following the mimicry, they were then asked to assess the emotional intensity of the seed video by selecting from a range of predefined emotional categories. The following emotion dimensions are targeted: ’Admiration’, ’Amusement’, ’Determination’, ’Empathic Pain’, ’Excitement’, and ’Joy’. A normalized score from 0 to 1 is provided as a ground truth value.
@@ 4.2 Expression Classification Challenge
baseline 0.25
@@ 4.4 Emotional Mimicry Intensity Estimation Challenge
We established baseline results using two different feature sets.
First, we employed pre-trained Vision Transformer (ViT) features, which were further processed through a three-layer Gated Recurrent Unit (GRU) network. This approach achieved a performance score of 0.09.
Secondly, we utilized features extracted from Wav2Vec2, combined with a linear processing layer, which resulted in a performance score of 0.24.
Additionally, we explored a multimodal approach by averaging the predictions from both unimodal methods, leading to a combined performance score of 0.48.
Table 6: Emotional Mimicry Intensity Estimation Challenge’s Results; the evaluation criterion is the average Pearson’s Correlation
Teams

### 10 Mar 2024 v2

@@ 1 Introduction
For this Challenge, an augmented version of the Aff-Wild2 [88, 36, 33, 38, 35, 34, 39, 31, 37, 41] is used. This corpora is audiovisual (A/V), in-the-wild and in total consists of 594 videos of around 3M frames of 584 subjects.
@@
Only uni-task solutions will be accepted for this Challenge. Teams are allowed to use any -publicly or not- available pre-trained model (as long as it has not been pre-trained on the utilized in this Challenge corpora, Aff-Wild2). The pre-trained model can be pre-trained on any task (e.g., VA estimation, Expression Recognition, AU detection, Face Recognition). However when the teams are refining the model and developing the methodology, they should not use any other annotations (VA or AUs): the methodology should be purely uni-task, using only the Expr annotations. This means that teams are allowed to use other databases’ Expr annotations, or generated/synthetic data (e.g. the data provided in the ECCV 2022 run of the ABAW Challenge [30]), or any affine transformations, or in general data augmentation techniques (e.g., [59]) for increasing the size of the training dataset.
@@
Only uni-task solutions will be accepted for this Challenge. Teams are allowed to use any -publicly or not- available pre-trained model (as long as it has not been pre-trained on the utilized in this Challenge corpora, Aff-Wild2). The pre-trained model can be pre-trained on any task (e.g., VA estimation, Expression Classification, AU detection, Face Recognition). However when the teams are refining the model and developing the methodology, they should not use any other annotations (VA or Expr): the methodology should be purely uni-task, using only the AU annotations. This means that teams are allowed to use other databases’ AU annotations, or generated/synthetic data, or any affine transformations, or in general data augmentation techniques (e.g., [59]) for increasing the size of the training dataset.
@@
The final fourth challenge is the Emotional Mimicry Intensity (EMI) Estimation challenge where emotional mimics are explored. Participants are asked to predict six emotional dimensions using a multi-output regression approach. The following emotional expressions have been used: ”Admiration”, ”Amusement”, ”Determination”, ”Empathic Pain”, ”Excitement”, and ”Joy”.
For the purposes of this challenge, we use the audiovisual and in-the-wild HUME-Vidmimic2 dataset, a comprehensive collection derived from ’in-the-wild’ settings, which contains more than 17 000 videos totaling over 30 hours from the United States, similar to our first version [7].
The sixth ABAW Competition, which is part of the respective Workshop to be held in conjunction with the IEEE Computer Vision and Pattern Recognition Conference (CVPR) 2024 is a continuation of the successful series of ABAW Competitions held in conjunction with IEEE CVPR 2023, ECCV 2022, IEEE CVPR 2022, ICCV 2021, IEEE FG 2020 and IEEE CVPR 2017, with the participation of many teams coming from both academia and industry, from all across the world [97, 92, 80, 64, 74, 79, 84, 99, 40, 98, 97, 64, 83, 87, 18, 54, 52, 99, 65, 29, 11, 46, 96, 13, 6, 4, 12, 94, 73, 75, 90, 81, 26, 1, 56, 43, 15, 14, 85, 49, 16, 50, 57, 24, 20, 10, 61, 51, 91, 55, 62, 27, 60, 95, 86, 28, 58, 82, 22, 78, 21, 67, 77, 25, 9, 19, 66, 5, 17, 47, 76, 53, 63, 48, 93, 44, 23, 45, 42].
@@ 2.5 Emotional Mimicry Intensity Estimation Challenge
In the Emotional Mimicry Intensity Challenge (EMI-Challenge), we investigate the study of emotional mimicry by presenting a large-scale and in-the-wild dataset featuring 557 participants and over 30 hours of audiovisual content. This dataset was collected in naturalistic settings, with participants using their webcams to record their facial and vocal responses by mimicking a ”seed” video and rating it in the range from 0 to 100.
The data preparation process involved a speaker-independent partitioning of the dataset into training, validation, and test sets. Table 3 statistics of the dataset for each partition. The train and validation data along with their corresponding annotations are provided to the participating teams. The unlabeled test data are provided to the participating teams who will upload their test set predictions to an evaluation server.
Along with the data, the participants are provided the faces of individuals within the videos that were detected with the use of MTCNN [89] at a frequency of 6 frames per second. In addition, features extracted from the raw signals and thus enabling participants to use end-to-end approaches [68, 71, 69, 72, 70] are provided. Specifically, the feature sets provided are the Vision Transformer (ViT) [3] for the faces and Wav2Vec 2.0 [2] for the audio signals.
Partition Duration # Samples
(HH:MM:SS)
Train 15:07:03 8072
Validation 9:12:02 4588
Test 9:04:05 4586
Table 3: HUME-Vidmimic2 partition statistics.
@@ 4.2 Expression Classification Challenge
baseline without MixAugment 0.23
baseline with MixAugment 0.25
@@ 4.3 Action Unit Detection Challenge
4.4 Compound Expression Recognition Challenge
No baseline network and results will be provided for this Challenge due to the nature of the Challenge.
@@ 4.5 Emotional Mimicry Intensity Estimation Challenge
We set initial baseline results with two distinct sets of features. Initially, we utilized features derived from a pre-trained Vision Transformer (ViT), which were then processed by a three-layer Gated Recurrent Unit (GRU) network, achieving a performance score of $\mathcal{P}_{EMI}=0.09$. Subsequently, we leveraged features extracted from Wav2Vec2, paired with a linear processing layer, attaining a performance score of $\mathcal{P}_{EMI}=0.24$. Furthermore, by averaging the predictions from both of these unimodal techniques, we pursued a multimodal strategy that culminated in an enhanced combined performance score of $\mathcal{P}_{EMI}=0.48$.
Table 7: Baseline results for the EMI challenge.
Model